OpenStudy (anonymous):

Factor this trig. expression: (sinx+1)² - (sinx-1)² Show steps. - is there a way of factoring without knowing a2−b2=(a−b)(a+b) ?

OpenStudy (ash2326):

It's of the form \[a^2-b^2=(a-b)(a+b)\] so using this here \[(\sin x+1)² - (\sin x-1)²= (\sin x+1+\sin x-1)(\sin x+1-(\sin x-1))\] Can you simplify the rest?

OpenStudy (anonymous):

\[(a+b)^2 - (a-b)^2 = 4ab\] a = sinx and b = 1 Put and solve...

OpenStudy (ganpat):

a2−b2=(a−b)(a+b) a = Sinx + 1, b = sinx -1 SO, [ (sinx +1) - (sinx -1)] [(sinx +1) +(sinx -1)] [sinx +1 - sinx +1] [sinx +1 + sinx - 1] [2] [2 sinx] So answer would be = 4 sinx ....